OpenStudy (anonymous):

The altitude of a triangle is increasing at a rate of 1.5 centimeters/minute while the area of the triangle is increasing at a rate of 1.0 square centimeters/minute. At what rate is the base of the triangle changing when the altitude is 9.5 centimeters and the area is 97 square centimeters?

OpenStudy (anonymous):

A = 1/2 b * h Using the product rule we get: dA/dt = 1/2 dh/dt * b + 1/2 db/dt * h We know the following dA/dt = 5 dh/dt = 1.5 db/dt = x (the thing we want to solve) h = 11.5 b = ? (we need to find this first) We know that A = (1/2)*b*h 97 = (1/2)*b*(11.5) we can now solve for b 97*2 = b * (23/2) b = (97*2*2) / 23 b = 388 / 23 Now we can solve for db/dt. dA/dt = 1/2 dh/dt * b + 1/2 db/dt * h 5 = (1/2)*(1.5)*(388/23) + (1/2)*(x)*(11.5) 5 = 291 / 23 + 23x / 4 5 - 291/23 = 23x / 4 -176 / 23 = 23x / 4 -176/23 * 4/23 = x x = -704 / 529 x = -1.33 cm/min